# ScaffoldGPT: A Scaffold-based GPT Model for Drug Optimization

### v2

@@ Abstract
Drug optimization has become increasingly crucial in light of fast-mutating virus strains and drug-resistant cancer cells. Nevertheless, it remains challenging as it necessitates retaining the beneficial properties of the original drug while simultaneously enhancing desired attributes beyond its scope. In this work, we aim to tackle this challenge by introducing ScaffoldGPT, a novel Generative Pretrained Transformer (GPT) designed for drug optimization based on molecular scaffolds. Our work comprises three key components: (1) A three-stage drug optimization approach that integrates pretraining, finetuning, and decoding optimization. (2) A uniquely designed two-phase incremental training approach for pre-training the drug optimization GPT on molecule scaffold with enhanced performance. (3) A token-level decoding optimization strategy, Top-N, that enabling controlled, reward-guided generation using pretrained/finetuned GPT. We demonstrate via a comprehensive evaluation on COVID and cancer benchmarks that ScaffoldGPT outperforms the competing baselines in drug optimization benchmarks, while excelling in preserving original functional scaffold and enhancing desired properties.
@@ 1 Introduction
On the other hand, generative pretrained language models have shown exceptional performance in diverse areas, from natural language understanding and generation exemplified by ChatGPT (Wu et al., 2023, Ouyang et al., 2022), to text-to-video conversion as seen in SORA (Liu et al., 2024c), and in programming and code generation through platforms like PG-TD (Zhang et al., 2023a) and Copilot (Nguyen and Nadi, 2022). Nevertheless, in the field of drug discovery, tools such as DrugGPT (Li et al., 2023), ChatDrug (Liu et al., 2024a) and ChemGPT (Frey et al., 2023), despite making some initial strides, have not yet achieved performance on par with their counterparts in other domains. The field of drug discovery remains in expectation of breakthroughs similar to the ones achieved by LMs in other domains.
Several challenges have hindered the impact of Language Models (LMs) on drug design: Firstly, the molecules created need to satisfy multiple criteria such as solubility and synthesizability, and must also secure a high docking score against a specific target site. However, existing drug discovery LMs generally only undergo pretraining with molecules and do not focus on enhancing multiple attributes simultaneously. Secondly, given that drugs sharing similar chemical structures tend to exhibit comparable biological or chemical effects (Bender and Glen, 2004), it is crucial to optimize the drug while preserving the beneficial chemical structure of the original molecule. Lastly, the current methodologies in drug discovery LMs primarily focus on maximizing likelihood during the decoding phase, rather than customizing the optimization process to meet specific goals.
REINVENT 4 (He et al., 2021; 2022, Loeffler et al., 2024) achieves the state of art performance in LM-based drug optimization by using the Transformer model and conducting experiments with randomly selected ligand pairs that maintain constrained Tanimoto Similarity. While this method successfully produces ligands with high Tanimoto Similarity, it encounters difficulties in consistently improving the properties of the original drug and tends to simply maximize the likelihood of molecule generation.
@@
$\bullet$ A framework for drug optimization using GPT that involves a three-step optimization process. The underlying motivation is that each stage is complementary to each other, enhances the performance of the preceding one. Furthermore, we have conducted ablation studies to illustrate the performance gains achieved at each step.
$\bullet$ A scaffolds-based GPT with a novel two-phase incremental training specifically designed for drug optimization. The motivation for incremental training is to conduct local optimization before embarking on global optimization by breaking down the whole training corpus into knowledge pieces in an incremental order. We also conducted an ablation study in table 3 to demonstrate the effectiveness of the incremental training.
@@ 2.1 Reinforcement Learning and Finetuning
Reinforcement learning (Tan et al., 2022a) has become a fundamental strategy in drug design (Born et al., 2021, Guimaraes et al., 2017, Neil et al., 2018, Olivecrona et al., 2017, Popova et al., 2018, Ståhl et al., 2019, Tan et al., 2022b, Wang et al., 2022a, Zhang et al., 2023b, Zhou et al., 2019), focusing on optimizing rewards that aggregate predicted property scores from various pharmaceutical predictors. Traditional reinforcement learning methods in drug discovery often neglected molecular structure constraints, leading to significant structural changes and unsynthesizable compounds. Our research differs by refining existing drugs to enhance their attributes without redesigning them from scratch and using reinforcement learning to improve a pre-trained language model generator, rather than starting a new. We also incorporate methods like Advantage-aligned Policy Optimization (APO) (Liu et al., 2023a), which assigns rewards based on advantage preference over the original molecule, to fine-tune a Transformer model, ensuring it aligns with multiple pharmaceutical objectives while preserving molecular structure. This approach, which includes controllable decoding, refines the model beyond traditional reinforcement learning fine-tuning stages.
@@ 2.2 Planning with GPT Models
Several studies have leveraged planning algorithms to enhance text outputs for a variety of NLP tasks. These include approaches like beam search optimization, machine translation improvements (Scialom et al., 2021, Leblond et al., 2021, Chaffin et al., 2021). The PG-TD (Zhang et al., 2023a) method is tailored for code generation using a singular reward function, whereas ERP (Liu et al., 2024b) introduces a novel concept by considering the certainty of each generated token along with an e-step forward entropy measurement to gauge potential outcomes. It has been shown that ERP effectively balances exploration and exploitation within molecular structures, leading to the discovery of high-reward molecules. Unlike previous studies that focus solely on planning with pre-trained language models, our approach incorporates a novel decoding optimization as a critical final step in the algorithm. Moreover, our focus is on optimizing an existing drug rather than creating a De Novo one from the ground up.
@@ 2.3 Drug Optimization
Recent drug optimization efforts have focused primarily on a limited array of drug properties while often disregarding the docking score, an essential metric for evaluating structural compatibility with a target (Zhou et al., 2019, Erikawa et al., 2021). DrugEx v3 (Liu et al., 2023b) seeks to resolve this deficiency by using 3D molecular graphs that encapsulate more comprehensive data such as chemical valence rules. Nevertheless, this method’s complexity poses challenges in generating molecules that closely resemble the originals. The resulting lower similarity could account for the diminished efficacy of graph-based methods, as deviations from the original molecular structures result in the loss of vital chemical properties. Conversely, our approach manages to preserve a decent level of similarity. Diffusion-based efforts (Alakhdar et al., 2024, Morehead and Cheng, 2024) either focus on De Novo drug discovery while neglecting the essential similarity to the original molecule, or overlook the docking score towards a binding target. Molsearch (Sun et al., 2022) is a Monte Carlo tree search (MCTS)-driven approach for molecular generation and optimization, and MIMOSA (Fu et al., 2021) is a GNN sampling-based method leveraging graph-based molecular optimization. DrugImprover (Liu et al., 2023a) effectively begins to redefine the drug optimization challenge by employing reinforcement learning with a mix of multiple objectives as rewards. Still, it employs an LSTM in its generative model, which faces limitations in scalability, capacity, and understanding context. In contrast, our method applies a Transformer as the primary generative model, enhanced further with Advantage-aligned Policy Optimization (APO) for exploratory purposes and an optimized decoder for superior performance.
@@ 2.4 Language Models for Drug Discovery
Large language models like MolGPT (Bagal et al., 2021) and ChemGPT (Frey et al., 2023) have been utilized in molecule generation and drug discovery (Bagal et al., 2021, Rothchild et al., 2021, Wang et al., 2022b), using formats like SMILES (Weininger, 1988) and SELFIES (Krenn et al., 2020) to standardize molecular representation. These models have shown promise in drug design, outperforming traditional methods in some predictive capacities. However, as noted by  Murakumo et al. (2023), using pre-trained language models typically results in minor molecular modifications and serves primarily supportive roles in the design process. We adopt the SMILES format and combine the strengths of GPTs with reinforcement learning. Unlike previous efforts that relied solely on LLMs, our approach uses RL fine-tuning to align with multiple pharmaceutical objectives and employs controlled decoding to guide the GPT model in generating more effective molecular structures.
Further developments include transformer-based models like REINVENT 4  (He et al., 2021; 2022, Loeffler et al., 2024), which mostly focuses on pretraining. While pretraining helps in generating molecules similar to those in the training dataset, it also inherently restricts the exploration scope due to biases in the training data. REINVENT4 uses the original molecule as input, generates molecules that are very similar to the original ones. This likely contributes to its relatively poor performance in other metrics, as the generated molecules exhibit minimal changes in chemical properties. In contrast, our proposed method, which uses SMILES and scaffolds as prompts, achieves a well-balanced trade-off between diversity and similarity in the generated and original molecules, potentially leading to improved performance.
@@ 3 Preliminaries
[…]
@@
[…]
@@ 4 The ScaffoldGPT Algorithm
13$\triangleright$ /* Stage 3: Token-level Decoding Optimization */
14:Optimize the generation of $\pi_{{\theta}}$ via Top-N (12) decoding strategy.
@@
In this project, we employ pre-training to harness large quantities of unlabeled text to construct a basic foundation model of language understanding. This foundation model can subsequently be fine-tuned and tailored to meet various specialized goals. In this work, we propose a novel framework for pre-training a Transformer and linking scaffolds with complete molecules, based on a SMILES (Simplified Molecular Input Line Entry System) (Weininger, 1988) string representation of the molecule. Here, we define $\mathcal{S}$ as the scaffold space and $\mathcal{Y}$ as the target molecule space. With $\mathcal{P}=\left\{\left({s,y}\right)|s,y\in\mathcal{S}\times\mathcal{Y}\right\}$, we denote the set of scaffold and molecular pairs from $\mathcal{S}$ and $\mathcal{Y}$. In this notation, $s$ represents the scaffold of the target molecule, and $y$ is the corresponding target molecule. We initially pre-trained our tokenizer using the Byte Pair Encoding (BPE) method (Gage, 1994, Sennrich et al., 2015). Building on the pre-trained BPE tokenizer, we propose a two-phase incremental training approach, as illustrated in Fig. 1, to notably enhance the model’s ability to improve the validity of inferring the target molecule from its scaffold.
Incremental training.
The rationale for incremental training is to conduct local optimization before embarking on global optimization. Therefore, we divide our training into two phases. In the first phase, we focus on training a GPT exclusively for molecules using Causal Language Modeling (CLM). CLM utilizes an autoregressive method where the model is trained to predict the next token in a sequence by considering only the tokens that precede it. The phase 1 corpus is designed as follows:
Figure 1: Two-phase incremental training of ScaffoldGPT. The first phase concentrates on recognizing the molecule, while the second phase builds connections between its scaffold and the original molecule.
$C_{\text{Phase 1}}=\left\{[BOS],\left\langle L\right\rangle,\underbrace{y_{1},\cdots,y_{T}}_{\text{target molecule Y}},[EOS]\right\},\left\langle L\right\rangle\text{is the token for ligand.}$. (4)
@@
$C_{\text{Phase 2}}=\left\{[BOS],\left\langle S\right\rangle,\underbrace{s_{1},\cdots,s_{T}}_{\text{source scaffold S}},\left\langle L\right\rangle,\underbrace{y_{1},\cdots,y_{T}}_{\text{target molecule Y}},[EOS]\right\},\left\langle S\right\rangle\text{: scaffold.}$. (5)
Consequently, the model can generate the appropriate molecule when given a scaffold. However, since a single scaffold may correspond to multiple molecules, we further refine the GPT-based generator policy, $\pi_{\theta}$, to target specific outcomes by applying reinforcement learning finetuning in the next stage.
Stage 2. RL finetuning. Fine-tuning a generative model is crucial for producing outcomes that meet specific objectives. In this study, we use the Advantage-alignment Policy Optimization (APO) (Liu et al., 2023a) algorithm to fine-tune the pretrained GPT-based generator policy $\pi_{\theta}$. This approach steers the model from a given scaffold towards the targeted molecule, simultaneously improving multiple properties.
In this work, we adopt the define of reward function from Liu et al. (2024b), and regarded each pharmaceutical property as a critic $C$ and got an ensemble critics $\mathbf{C}$ as follows:
@@
where $\text{where }Y_{j}=$ $\left[\mathbf{y}_{<i},y_{i},\cdots,y_{T}\right]_{j},\text{and }y_{i}\sim\textsc{Top-PK}\left({\mathbf{y}_{<i},p,k}\right)|_{S}.$
where Top-PK (Liu et al., 2024b) is defined as follows:
[…]
@@
[…]

### v3

@@ Abstract
Drug optimization has become increasingly crucial in light of fast-mutating virus strains and drug-resistant cancer cells. Nevertheless, it remains challenging as it necessitates retaining the beneficial properties of the original drug while simultaneously enhancing desired attributes beyond its scope. In this work, we aim to tackle this challenge by introducing ScaffoldGPT, a novel Generative Pretrained Transformer (GPT) designed for drug optimization based on molecular scaffolds. Our work comprises three key components: (1) A three-stage drug optimization approach that integrates pretraining, finetuning, and decoding optimization. (2) A novel two-phase incremental pre-training strategy for scaffold-based drug optimization. (3) A token-level decoding optimization strategy, Top-N, that enabling controlled, reward-guided generation using the pretrained or finetuned GPT. We demonstrate via a comprehensive evaluation on COVID and cancer benchmarks that ScaffoldGPT outperforms the competing baselines in drug optimization benchmarks, while excelling in preserving original functional scaffold and enhancing desired properties.
@@ 1 Introduction
On the other hand, generative pretrained language models have shown exceptional performance in diverse areas, from natural language understanding and generation exemplified by ChatGPT (Wu et al., 2023; Ouyang et al., 2022), to text-to-video conversion as seen in SORA (Liu et al., 2024d), and in programming and code generation through platforms like PG-TD (Zhang et al., 2023a) and Copilot (Nguyen and Nadi, 2022). Nevertheless, in the field of drug discovery, tools such as DrugGPT (Li et al., 2023), ChatDrug (Liu et al., 2024a) and ChemGPT (Frey et al., 2023), despite making some initial strides, have not yet achieved performance on par with their counterparts in other domains. The field of drug discovery remains in expectation of breakthroughs similar to the ones achieved by LMs in other domains.
Several challenges have hindered the impact of Language Models (LMs) on drug design: Firstly, the molecules created need to satisfy multiple criteria such as solubility and synthesizability, and must also secure a high docking score against a specific target site (Liu et al., 2024b). However, existing drug discovery LMs generally only undergo pretraining with molecules and do not focus on enhancing multiple attributes simultaneously. Secondly, given that drugs sharing similar chemical structures tend to exhibit comparable biological or chemical effects (Bender and Glen, 2004), it is crucial to optimize the drug while preserving the beneficial chemical structure of the original molecule. Lastly, the current methodologies in drug discovery LMs primarily focus on maximizing likelihood during the decoding phase, rather than customizing the optimization process to meet specific goals.
REINVENT 4 (He et al., 2021, 2022; Loeffler et al., 2024) achieves the state of art performance in LM-based drug optimization by using the Transformer model and conducting experiments with randomly selected ligand pairs that maintain constrained Tanimoto Similarity. While this method successfully produces ligands with high Tanimoto Similarity, it encounters difficulties in consistently improving the properties of the original drug and tends to simply maximize the likelihood of molecule generation.
Generalizable Insights about Machine Learning in the Context of Healthcare
@@
$\bullet$ A novel framework for drug optimization using GPT, featuring a three-step optimization process, the first of its kind to the best of our knowledge. The underlying motivation is that each stage is complementary to each other, enhances the performance of the preceding one. Furthermore, we have conducted ablation studies to illustrate the performance gains achieved at each step.
$\bullet$ A scaffolds-based GPT with a novel two-phase incremental pre-training specifically designed for drug optimization. The motivation for incremental training is to conduct local optimization before embarking on global optimization by breaking down the whole training corpus into knowledge pieces in an incremental order. We also conducted an ablation study in table 3 to demonstrate the effectiveness of the incremental training.
@@ 2.1 Reinforcement Learning and Finetuning
Reinforcement learning (Tan et al., 2022a; Liu et al., 2023c, b) has become a fundamental strategy in drug design (Born et al., 2021; Guimaraes et al., 2017; Neil et al., 2018; Olivecrona et al., 2017; Popova et al., 2018; Ståhl et al., 2019; Tan et al., 2022b; Wang et al., 2022a; Zhang et al., 2023b; Zhou et al., 2019), focusing on optimizing rewards that aggregate predicted property scores from various pharmaceutical predictors. Traditional reinforcement learning methods in drug discovery often neglected molecular structure constraints, leading to significant structural changes and unsynthesizable compounds. Our research differs by refining existing drugs to enhance their attributes without redesigning them from scratch and using reinforcement learning to improve a pre-trained language model generator, rather than starting a new. We also incorporate methods like Advantage-aligned Policy Optimization (APO) (Liu et al., 2023a), which assigns rewards based on advantage preference over the original molecule, to fine-tune a Transformer model, ensuring it aligns with multiple pharmaceutical objectives while preserving molecular structure. This approach, which includes controllable decoding, refines the model beyond traditional reinforcement learning fine-tuning stages.
@@ 2.2 Planning with GPT Models
Several studies have leveraged planning algorithms to enhance text outputs for a variety of NLP tasks. These include approaches like beam search optimization, machine translation improvements (Scialom et al., 2021; Leblond et al., 2021; Chaffin et al., 2021). The PG-TD (Zhang et al., 2023a) method is tailored for code generation using a singular reward function, whereas ERP (Liu et al., 2024c) introduces a novel concept by considering the certainty of each generated token along with an e-step forward entropy measurement to gauge potential outcomes. It has been shown that ERP effectively balances exploration and exploitation within molecular structures, leading to the discovery of high-reward molecules. Unlike previous studies that focus solely on planning with pre-trained language models, our approach incorporates a novel decoding optimization as a critical final step in the algorithm. Moreover, our focus is on optimizing an existing drug rather than creating a De Novo one from the ground up.
@@ 2.3 Drug Optimization
Recent drug optimization efforts have focused primarily on a limited array of drug properties while often disregarding the docking score, an essential metric for evaluating structural compatibility with a target (Zhou et al., 2019; Erikawa et al., 2021). DrugEx v3 (Liu et al., 2023d) seeks to resolve this deficiency by using 3D molecular graphs that encapsulate more comprehensive data such as chemical valence rules. Nevertheless, this method’s complexity poses challenges in generating molecules that closely resemble the originals. The resulting lower similarity could account for the diminished efficacy of graph-based methods, as deviations from the original molecular structures result in the loss of vital chemical properties. Conversely, our approach manages to preserve a decent level of similarity.
Diffusion-based efforts (Alakhdar et al., 2024; Morehead and Cheng, 2024) either focus on De Novo drug discovery while neglecting the essential similarity to the original molecule, or overlook the docking score towards a binding target. Molsearch (Sun et al., 2022) is a Monte Carlo tree search (MCTS)-driven approach for molecular generation and optimization, and MIMOSA (Fu et al., 2021) is a GNN sampling-based method leveraging graph-based molecular optimization. DrugImprover (Liu et al., 2023a) effectively begins to redefine the drug optimization challenge by employing reinforcement learning with a mix of multiple objectives as rewards. Still, it employs an LSTM in its generative model, which faces limitations in scalability, capacity, and understanding context. In contrast, our method applies a Transformer as the primary generative model, enhanced further with Advantage-aligned Policy Optimization (APO) for exploratory purposes and an optimized decoder for superior performance.
@@ 2.4 Language Models for Drug Discovery
Large language models like MolGPT (Bagal et al., 2021) and ChemGPT (Frey et al., 2023) have been utilized in molecule generation and drug discovery (Bagal et al., 2021; Rothchild et al., 2021; Wang et al., 2022b), using formats like SMILES (Weininger, 1988) and SELFIES (Krenn et al., 2020) to standardize molecular representation. These models have shown promise in drug design, outperforming traditional methods in some predictive capacities. However, as noted by  Murakumo et al. (2023), using pre-trained language models typically results in minor molecular modifications and serves primarily supportive roles in the design process. We adopt the SMILES format and combine the strengths of GPTs with reinforcement learning. Unlike previous efforts that relied solely on LLMs, our approach uses RL fine-tuning to align with multiple pharmaceutical objectives and employs controlled decoding to guide the GPT model in generating more effective molecular structures.
Further developments include transformer-based models like REINVENT 4  (He et al., 2021, 2022; Loeffler et al., 2024), which mostly focuses on pretraining. While pretraining helps in generating molecules similar to those in the training dataset, it also inherently restricts the exploration scope due to biases in the training data. REINVENT4 uses the original molecule as input, generates molecules that are very similar to the original ones. This likely contributes to its relatively poor performance in other metrics, as the generated molecules exhibit minimal changes in chemical properties. In contrast, our proposed method, which uses SMILES and scaffolds as prompts, achieves a well-balanced trade-off between diversity and similarity in the generated and original molecules, potentially leading to improved performance.
@@ 3 Preliminaries
LM. We define the state space $\mathcal{S}$ as the set of all possible molecule, where each molecule is represented as a state $s$ that includes a start token $\left[\text{BOS}\right]$, a molecule with SMILES (Weininger, 1988) representation string, and a termination action $\left[\text{EOS}\right]$. We define the set of complete molecules as
@@
where $\mathcal{Y}_{t}\subseteq\mathcal{S}_{t}|_{t\in\left[T\right]}$ represents the hypothesis space at step $t$ (sequence length $t$), $\mathcal{V}^{*}$ represents the Kleene closure of Transformer’s vocabulary $\mathcal{V}$, with $\mathcal{V}:=\mathcal{A}$, and $\circ$ indicating string concatenation. Each action $a\in\mathcal{A}$ is represented as token $y\in\mathcal{V}$. In this work, we train a GPT policy $\pi_{\theta}$ to acquire prior knowledge for generating valid molecules based on a given set of molecules $\mathcal{B}$. We define the generator policy $\pi_{\theta}$, with learned weights $\theta$, as the product of probability distributions: $\pi_{\theta}\left({\mathbf{y}|\mathbf{x}}\right)=\prod_{t=1}^{|\mathbf{y}|}\pi_{\theta}\left({y_{t}|\mathbf{x},\mathbf{y}_{<t}}\right)$, where $\pi_{\theta}\left({\cdot|\mathbf{x},\mathbf{y}_{<t}}\right)$ is a distribution, $\mathbf{x}$ is an input sequence, and $\mathbf{y}_{<1}=y_{0}:=\text{[BOS]}$. The decoding process in text generation involves identifying the most likely hypothesis by optimizing the objective: $\mathbf{y}^{\star}=\mathop{\rm arg\,max}_{\mathbf{y}\in\mathcal{Y}_{T}}\log\pi_{\theta}\left({\mathbf{y}|\mathbf{x}}\right).$
Figure 1: Overview of ScaffoldGPT. (a) ScaffoldGPT decomposes molecular knowledge top-down into fragments and then reconstructs it bottom-up through linking. (b) In Stage 1, it proposes an incremental pretraining strategy. (c) In Stage 2, it undergoes reinforcement learning-based post-training. (d) In Stage 3, it introduces a token-level controllable decoding optimization approach. ScaffoldGPT is the first to introduce a three-stage optimization framework, where each stage complements the others.
@@ 4 The ScaffoldGPT Algorithm
13:end for
[…]
15:Optimize the generation of $\pi_{{\theta}}$ via Top-N (13) decoding strategy.
@@
In this project, we employ pre-training to harness large quantities of unlabeled text to construct a basic foundation model of language understanding. This foundation model can subsequently be fine-tuned and tailored to meet various specialized goals. In this work, we propose a novel framework for pre-training a Transformer and linking scaffolds with complete molecules, based on a SMILES (Simplified Molecular Input Line Entry System) (Weininger, 1988) string representation of the molecule. Here, we define $\mathcal{S}$ as the scaffold space and $\mathcal{Y}$ as the target molecule space. With $\mathcal{P}=\left\{\left({s,y}\right)|s,y\in\mathcal{S}\times\mathcal{Y}\right\}$, we denote the set of scaffold and molecular pairs from $\mathcal{S}$ and $\mathcal{Y}$. In this notation, $s$ represents the scaffold of the target molecule, and $y$ is the corresponding target molecule. We initially pre-trained our tokenizer using the Byte Pair Encoding (BPE) method (Gage, 1994; Sennrich et al., 2015). Building on the pre-trained BPE tokenizer, we propose a two-phase incremental training approach, as illustrated in Fig. 2, to notably enhance the model’s ability to improve the validity of inferring the target molecule from its scaffold.
Incremental pre-training.
The rationale for incremental pre-training is to conduct local optimization before embarking on global optimization. Therefore, we divide our training into two phases. In the first phase, we focus on training a GPT exclusively for molecules using Causal Language Modeling (CLM). CLM utilizes an autoregressive method where the model is trained to predict the next token in a sequence by considering only the tokens that precede it. The phase 1 corpus is designed as follows:
Figure 2: Two-phase incremental training of ScaffoldGPT. The first phase concentrates on recognizing the molecule, while the second phase builds connections between its scaffold and the original molecule. This process can be extended into a three-phase pretraining framework by establishing connections among the molecule’s fragments, scaffold, and original molecule.
$C_{\text{Phase 1}}=\left\{[BOS],\left\langle L\right\rangle,\underbrace{y_{1},\cdots,y_{T}}_{\text{target molecule Y}},[EOS]\right\},\left\langle L\right\rangle\text{is the token for ligand.}$ (4)
@@
$C_{\text{Phase 2}}=\left\{[BOS],\left\langle S\right\rangle,\underbrace{s_{1},\cdots,s_{T}}_{\text{source scaffold S}},\left\langle L\right\rangle,\underbrace{y_{1},\cdots,y_{T}}_{\text{target molecule Y}},[EOS]\right\},\left\langle S\right\rangle\text{: scaffold.}$ (5)
Consequently, the model can generate the appropriate molecule when given a scaffold as prompt. While this work primarily focuses on scaffolds, the proposed method can be extended to fragments as well. A scaffold represents a complete core structure, while fragments may constitute parts of the scaffold or side chains attached to it. This approach could be further developed into a fragment-based third-phase pre-training strategy, with the phase 3 corpus outlined as follows:
$C_{\text{Phase 3}}=\left\{[BOS],\left\langle F\right\rangle,\underbrace{f_{1},\cdots,f_{T}}_{\text{source fragment F}},\left\langle S\right\rangle,\underbrace{s_{1},\cdots,s_{T}}_{\text{source scaffold S}},[EOS]\right\},\left\langle F\right\rangle\text{: fragment.}$ (6)
As a result, the model can produce a suitable scaffold when provided with a fragment as a prompt, and using the scaffold as the input prompt will generate a complete molecule. However, since a single scaffold can relate to multiple molecules and a fragment can correspond to multiple scaffolds, we further optimize the GPT-based generator policy, $\pi_{\theta}$, to focus on specific outcomes by implementing reinforcement learning fine-tuning in the subsequent stage.
Stage 2. RL post-training/finetuning. Fine-tuning a generative model is crucial for producing outcomes that meet specific objectives. In this study, we use the Advantage-alignment Policy Optimization (APO) (Liu et al., 2023a) algorithm to fine-tune the pretrained GPT-based generator policy $\pi_{\theta}$. This approach steers the model from a given scaffold towards the targeted molecule, simultaneously improving multiple properties.
In this work, we adopt the define of reward function from Liu et al. (2024c), and regarded each pharmaceutical property as a critic $C$ and got an ensemble critics $\mathbf{C}$ as follows:
@@
$\text{where }Y_{j}=$ $\left[\mathbf{y}_{<i},y_{i},\cdots,y_{T}\right]_{j},\text{and }y_{i}\sim\textsc{Top-PK}\left({\mathbf{y}_{<i},p,k}\right)|_{S}.$
where Top-PK (Liu et al., 2024c) is defined as follows:
$\text{{Top-PK}}\left({\mathbf{y}_{<i},p,k}\right)|_{S}=\mathcal{A}_{\mathbf{y}_{<i}},\text{where~}\mathcal{A}_{\mathbf{y}_{<i}}=\left\{y_{1},\ldots,y_{j}\right\},y_{i}\in\mathcal{V},$ (9)
@@
$Y^{\star}$ $\sim\left\{[\mathbf{y}_{<i},y_{i},\cdots,y_{T}]\right\},\text{~where~}y_{i}\sim\text{{Top-N}}\left({\mathbf{y}_{<i},p,k,n}\right)|_{S}$ (13)
 $\text{~{Top-N}}\left({\mathbf{y}_{<i},p,k,n}\right)|_{S}=\mathcal{A}_{\mathbf{y}_{<i}},\text{~}\mathcal{A}_{\mathbf{y}_{<i}}=\left\{y_{1},\ldots,y_{n}\right\},y_{i}\in\mathcal{V},|\mathcal{A}|\leq k,$